OpenStudy (mayankdevnani):

Find the remainder when 5^{2009}+13^{2009} is divided by 18.

OpenStudy (mayankdevnani):

@amistre64

OpenStudy (mayankdevnani):

@phi and @kelliegirl33

OpenStudy (phi):

I would look for a pattern of remainders when you divide 18 into 5, 5^1, 5^2, ... eventually it will start repeating. then reduce 5^2009 using modulo

OpenStudy (mayankdevnani):

i don't understand

OpenStudy (phi):

can you list the remainders you get when you divide 18 into 5^0, 5^1, 5^2, 5^3, 5^4, 5^5, 5^6, 5^7 ?

OpenStudy (mayankdevnani):

can you give me shortest method to solve it? it seems to be longer

OpenStudy (phi):

if you know a faster way, what is it ?

OpenStudy (mayankdevnani):

i don't know

OpenStudy (amistre64):

mod 18 would have been my suggestions as well.

OpenStudy (phi):

5^0, 5^1, 5^2, 5^3, 5^4, 5^5, 5^6, 5^7 ? rem: 1 5 7 17 13 11 1 5 it repeats modulo 6 for example 5^7/18 is the same as 5^(7 mod 6)/18 = 5/18 = 5 that means 5^2009 / 18 will have the same remainder as 5^n /18 where n = 2009 mod 6 = 5 looking at the table, we see 5^5/18 has a remainder of 11 now do the same thing for 13 add the two numbers, and find the remainder after dividing by 18

ganeshie8 (ganeshie8):

another quick way is to notice that, 'leaving 13 as remainder' is same as 'leaving -5 as remainder -5', so both terms wud simply cancel out giving u 0 : 5^{2009}+13^{2009} 5^{2009} - 5^{2009}

OpenStudy (phi):

ooh, that is quicker. @ganeshie8

OpenStudy (raden):

alternative : a^n + b^n = (a+b)(bla bla bla bla) for a,b are odd numbers so, 5^{2009}+13^{2009 = (5+13)(.........) = 18(.......) obvious, if it divided by 18, then its remainder is zero

OpenStudy (phi):

yes, that is a good way too. @RadEn

OpenStudy (mayankdevnani):

@ganeshie8 can you explain your answer please ?

OpenStudy (mayankdevnani):

@RadEn if a,b are even no.then

OpenStudy (mayankdevnani):

@phi i don't understand your answer

OpenStudy (mayankdevnani):

@amistre64 , @UnkleRhaukus @mukushla

OpenStudy (mayankdevnani):

and i don't understand "mod" also ?What is the significance of this ?

OpenStudy (raden):

yeah, it just for a,b are odd's

OpenStudy (mayankdevnani):

if a,b are even, is there any shortcut method?

OpenStudy (raden):

i never see for a,b are even

OpenStudy (mayankdevnani):

ohkk

OpenStudy (mayankdevnani):

can you tell me what is "mod"?

ganeshie8 (ganeshie8):

\(\large 5^{2009}+13^{2009}\) first observation : if \(a\) leaves a remainder 'r' when divided by some number, then \(a^n\) leaves a remainder of \(r^n\) next : since \(13\) leaves a remainder of \(-5\) when divided by 18, \(13^{2009}\) leaves a remainder of \((-5)^{2009}\) plugging them cancels out the left hand side completely giving u 0

ganeshie8 (ganeshie8):

mod is just a notation referring to remainders : \(a \equiv b \mod n\) means \(a-b\) is divisible by \(n\)

OpenStudy (mayankdevnani):

-5? how?

OpenStudy (mayankdevnani):

not +5

ganeshie8 (ganeshie8):

13/18 = (18*1 - 5) / 18 = 1 R (-5) = 0 R (13)

ganeshie8 (ganeshie8):

leaving 13 remainder is same as saying leaving 13-18 = -5 remainder

OpenStudy (mayankdevnani):

okk

ganeshie8 (ganeshie8):

basically, we can have both negative and positive remainders

OpenStudy (mayankdevnani):

how?

ganeshie8 (ganeshie8):

2/3 = 0 R (2) = 1 R (-1)

ganeshie8 (ganeshie8):

when u divide 3 in 2, u can ask two questions : 1) how much im overflowing 2) how much im lacking

ganeshie8 (ganeshie8):

answer to first question gives u positive remainder answer to second question gives u negative remainder

ganeshie8 (ganeshie8):

both are correct in remainder arithmetic

OpenStudy (mayankdevnani):

so then, what you want to say ?

ganeshie8 (ganeshie8):

wat do u mean ?

OpenStudy (mayankdevnani):

i mean that the remainder is not +ve in this case,right?

ganeshie8 (ganeshie8):

u get to choose, if u wanto use positive / negative remainder when u talk about dividing wid 18 : 13 = -5 = 31 = 13 + 18k

ganeshie8 (ganeshie8):

more precicely : \(13 \equiv -5 \equiv 31 \mod 18\)

ganeshie8 (ganeshie8):

forget it, if that doesnt make sense... just see this :- 2/3 = 0 R (2) = 1 R (-1)

ganeshie8 (ganeshie8):